OpenStudy (anonymous):

h(x) = 1 / (1+3t^2) Give the eighth Taylor polynomial of h about 0. Is there a way to do this without taking derivatives 8 times?

OpenStudy (inkyvoyd):

Find the rule.

OpenStudy (anonymous):

1-3x^2+9x^4-27x^6+-.....+(-1)^n.3^2n.x^2n+...

OpenStudy (anonymous):

1+q+q2+q3....=1/(1-q) this is rule

OpenStudy (anonymous):

q=-3x2

OpenStudy (anonymous):

@mahmit2012 How did you find the rule? I don't get what you did in your first post, did you plug in values?

OpenStudy (anonymous):

you can calculate 1+q+q2+.....q^(N-1)=(1-q^N)/1-q if abs(q)<1 and N go to inf you get my simply rule. have fun.!!